# A framework for the use of generative modelling in non-equilibrium statistical mechanics

### v3

@@ Abstract
We discuss an approach to mathematically modelling systems made of objects that are coupled together, using generative models of the dependence relationships between states (or trajectories) of the things comprising such systems. This broad class includes open or non-equilibrium systems and is especially relevant to self-organising systems. The ensuing variational free energy principle (FEP) has certain advantages over using random dynamical systems explicitly, notably, by being more tractable and offering a parsimonious explanation of why the joint system evolves in the way that it does, based on the properties of the coupling between system components. Using the FEP allows us to model the dynamics of an object as if it were a process of variational inference, because variational free energy (or surprisal) is a Lyapunov function for its dynamics. In short, we argue that using generative models to represent and track relations among subsystems leads us to a particular statistical theory of interacting systems. Conversely, this theory enables us to construct nested models that respect the known relations among subsystems. We point out that the fact that a physical object conforms to the FEP does not necessarily imply that this object performs inference in the literal sense; rather, it is a useful explanatory fiction which replaces the ‘explicit’ dynamics of the object with an ‘implicit’ flow on free energy gradients—a fiction that may or may not be entertained by the object itself.
@@ Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the third named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed. We thank the organisers and audiences of both. DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York. KF and DARS acknowledge support from the VERSES Research Lab.
@@ 1 Introduction
One of the points this paper will emphasise is that using this framework does not necessarily presuppose that a system actually performs inference on its environment. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. We need not assume that the particle itself is literally performing inference when we build models under our framework. What is at stake is an ‘as if’ description: we simply assume that there exists a quantity that varies systematically with the dynamics of the system; it so happens that surprisal (or, equivalently, variational free energy) is a good choice of quantity for physical reasons, and that the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator. The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane. It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems. However, it is not necessarily a faithful representation of the system in the literalist sense. See also [6].
@@ 2.1 Fundamentals
The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where ^$\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3]. Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution. In particular: if we assume that $\hat{\eta}_{b}$ parametrises $p(\eta\mid b)$, then we have made $\mu$ into an estimator of the statistics of the environment.
The FEP is simplest when we have a non-equilibrium steady state density, in which case the following claim holds: if the system remains on some attractor of states, it looks like it minimises variational free energy. This is true in the following sense. It is possible to express the dynamics of a random variable with a non-equilibrium steady state density in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
@@
Now observe that since the divergence term is non-negative, $F(\mu,b)$ is an upper bound on the surprisal $-\log p(\mu,b)$,
@@
Our reasoning so far implies this suffices as an approximate description of all Markov blanketed systems.
@@ 2.2 Analysis and interpretation
2.3 Empirical validation
In this section we bolster the claim that modelling physical systems with variational free energy is advantageous by reviewing some simulations self-organisation and self-assembly—namely, that of cellular morphogenesis. In these simulations we will see that introducing the generative model allows us to model the system self-organising as if it ‘knew its place’ within a larger system—inferring the unknowable external states by the minimisation of variational free energy. [^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/). These simulations can be reproduced using the morphogenesis feature and the data provided there.
Starting from an ensemble of undifferentiated cells at the same location, we wanted to simulate their migration and differentiation using the minimisation of variational free energy. The simulation here exploits the fact one can express a Markov blanketed system’s marginal Lyapunov function in terms of a variational free energy function, such that its attracting set can be prescribed in terms of the generative model that defines the free energy. In essence, the systems dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences.
A group of eight cells was simulated with states consisting of the physical locations on a two-dimensional grid ($\psi_{x}$) of each cell and the cell type of each cell (measured by its cell signalling profile, a four-dimensional vector of signals expressed $\psi_{c}$). For cell $i$, $(\psi_{xi},\psi_{ci})$ is an internal state, and the remaining tuples $(\psi_{xj},\psi_{cj})$ constitute external states.
We assume the active states of each cell couple to $\psi_{c}$ so that the actions the cell can take are manipulation of the signals expressed in the environment. We moreover assume the active states fluctuate quickly enough that from the viewpoint of the environment their dynamics are negligible, i.e. $\psi\approx a$, constituting an adiabatic approximation
@@ 2.3 Empirical validation
with large time constant $\tau$. Hence the $\psi_{xj}$’s become external states coupled to active states in the Markov blankets of the other cells.
@@
A point in the state space $(\psi_{x}^{*},\psi_{c}^{*})$ was designated as a target morphology, giving prescribed locations and cell types of all cells. This point consisted of the physical form of a head, body, and tail, and with different cell types in each component of the organism. We then equipped each cell with a generative model, specifying it as follows.
The mean was given in terms of how environmental states are mapped to sensory states,
@@
with the softmax function returning the expected identity of each cell. A precision $\Pi^{(2)}=e^{-2}$ was also set over external states. A prediction error for how the generative model matched with sensations was given in the form
@@
The generative model was operationalised by following the procedure in [14]. A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time. A derivative operator $D$ was defined for ordinary differential equations on this space. The entire trajectory was given a coordinate in function space
@@
Treating (2.1.3) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final equations of motion given the generative model were obtained as
$\begin{split}f_{\tilde{\psi}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})&=(Q_{\psi_{i}}-\Gamma_{\psi_{i}})\nabla_{\tilde{\psi}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})=D\tilde{\psi}_{i}-\nabla_{\tilde{\psi}_{i}}\tilde{\varepsilon}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}-\Pi^{(2)}\tilde{\psi}_{i}\\ f_{\tilde{a}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})&=(Q_{a_{i}}-\Gamma_{a_{i}})\nabla_{\tilde{a}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})=D\tilde{a}_{i}-\nabla_{\tilde{a}_{i}}\tilde{s}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}\\ \partial_{t}\tilde{a}_{c}&=D\tilde{a}_{c}-\Pi^{(1)}_{c}\tilde{\varepsilon}_{c}+\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}\\ \partial_{t}\tilde{a}_{x}&=D\tilde{a}_{x}-\nabla_{x}\tilde{s}_{x}\cdot\Pi^{(1)}_{x}\tilde{\varepsilon}_{x}+\nabla_{x}\tilde{s}_{\lambda}\cdot\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}.\end{split}$ (2.3.1)
We then simulated the dynamics of each cell.
@@
Figure 1: Plots of the sensed extracellular gradients (1(a)), encoded signal in the cells (1(b)) and softmax expectations of the identities of each cell (1(c)). In Figure 1(a) colours denote different signal expressions and black dots denote no data, with the final locations of the cells starred.
@@
Figure 2: Here morphogenesis is demonstrated by cells migrating to align with sensory input, modelled by a concomitant decrease of free energy in time.
@@
Figure 3: Dynamics of each cell, with final locations starred. The gradients here match those of Figure 2(a). It can be seen that (speaking somewhat heuristically) the cell moves to fulfil its expectations about the signals it will encounter, whilst expressing the signals associated with its current beliefs about its place in the target ensemble.

### v4

@@ Abstract
We discuss an approach to mathematically modelling systems made of objects that are coupled together, using generative models of the dependence relationships between states (or trajectories) of the things comprising such systems. This broad class includes open or non-equilibrium systems and is especially relevant to self-organising systems. The ensuing variational free energy principle (FEP) has certain advantages over using random dynamical systems explicitly, notably, by being more tractable and offering a parsimonious explanation of why the joint system evolves in the way that it does, based on the properties of the coupling between system components. The FEP is a method whose use allows us to build a model of the dynamics of an object as if it were a process of variational inference, because variational free energy (or surprisal) is a Lyapunov function for its dynamics. In short, we argue that using generative models to represent and track relations amongst subsystems leads us to a particular statistical theory of interacting systems. Conversely, this theory enables us to construct nested models that respect the known relations amongst subsystems. We point out that the fact that a physical object conforms to the FEP does not necessarily imply that this object performs inference in the literal sense; rather, it is a useful explanatory fiction which replaces the ‘explicit’ dynamics of the object with an ‘implicit’ flow on free energy gradients—a fiction that may or may not be entertained by the object itself.
@@ Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the third named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed. We thank the organisers and audiences of both. DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York. KJF and DARS acknowledge support from the VERSES Research Lab.
@@
Author contributions
The list of authors has been ordered alphabetically by surname. The CRediT contributions are as follows: KJF (formal analysis, software, writing—review and editing), MJDR (conceptualisation, writing—review and editing), DARS (conceptualisation, formal analysis, software, visualisation, writing—original draft).
@@ 1 Introduction
One of the points this paper will emphasise is that using this framework does not necessarily presuppose that a system actually performs inference on its environment. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. We need not assume that the particle itself is literally performing inference when we build models under our framework. What is at stake is an ‘as if’ description: we simply assume that there exists a quantity that varies systematically with the dynamics of the system; it so happens that surprisal (or, equivalently, variational free energy) is a good choice of quantity for physical reasons, and that the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator. The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane. It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems. However, it is not necessarily a faithful representation of the system in the literalist sense. (See also [6].)
@@ 2.1 Fundamentals
The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where $\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3]. Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution. In particular: if we assume that $\hat{\eta}_{b}$ parametrises $p(\eta\mid b)$, then we have made $\mu$ into an estimator of the statistics of the environment.
The FEP is simplest when we have a non-equilibrium steady state density, in which case the following claim holds: if the system remains on some attractor of states, it looks like it minimises variational free energy. This is true in the following sense. It is possible to express the dynamics of a random variable with a non-equilibrium steady state density
[…]
in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
[…]
where $Q(x)$ is skew-symmetric everywhere, $\Gamma(x)$ is positive (semi-)definite everywhere and satisfies $2\Gamma(x)=D(x)D^{\top}(x)$, and $p^{*}(x)$ is the non-equilibrium steady state density. This decomposition holds whenever there exists a $Q(x)$ such that
$\nabla\times[Q(x)+\Gamma(x)]^{-1}f(x)=0$
@@
Now observe that since the divergence term is non-negative, $F(\mu,b)$ is an upper bound on the surprisal $\log p(\mu,b)$,
@@
Evidently, when the inference gap is minimised, we have
[…]
on the nose. Our reasoning so far implies this suffices as an approximate description of all Markov blanketed systems.
@@ 2.2 Analysis and interpretation
3 Empirical validation
In this section we bolster the claim that modelling physical systems with variational free energy is advantageous by reviewing some simulations of self-organisation and self-assembly. Two toy models will be studied—cellular morphogenesis, and an ensemble of excitable cells. In these simulations we will see that introducing the generative model allows us to model the system self-organising as if it ‘knew its place’ within a larger system—inferring the unknowable external states by the minimisation of variational free energy.
3.1 Cellular morphogenesis
Starting from an ensemble of undifferentiated cells at the same location, we wanted to simulate their migration and differentiation using the minimisation of variational free energy. The simulation here exploits the fact one can express a Markov blanketed system’s marginal Lyapunov function in terms of a variational free energy function, such that its attracting set can be prescribed in terms of the generative model that defines the free energy. In essence, the systems dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences. [^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/). They can be reproduced using the morphogenesis feature and the data provided there.
A group of eight cells was simulated with states consisting of the physical locations on a two-dimensional grid ($\psi_{x}$) of each cell and the cell type of each cell (measured by its cell signalling profile, a four-dimensional vector of signals expressed $\psi_{c}$). For cell $i$, $(\psi_{xi},\psi_{ci})$ is an internal state $\mu$, and the remaining tuples $(\psi_{xj},\psi_{cj})$ constitute external states $\eta$.
Each cell was equipped with active states $(a_{x},a_{c})$ corresponding to modulations of its position and chemotactic signalling profile, respectively. We assume the active states fluctuate quickly enough that from the viewpoint of the environment their dynamics are negligible, i.e. $\psi\approx a$, constituting an adiabatic approximation
@@ 3.1 Cellular morphogenesis
with large time constant $\tau$. In this way the cell’s active states become the external states offered to other cells. (Note that this assumption is purely for convenience, simplifying the analytic expressions for certain quantities; it is by no means necessary in the following discussion, and the example to follow will not use it. What is at stake in terms of the realism of the simulation is: the adiabatic assumption excludes the possibility of modelling cell migration or signal expression as a function of action.)
@@
In summary, the active states are (i) where a cell chooses to be and (ii) what a cell chooses to release. Those actions are the external states being offered to other cells under the adiabatic assumption, and through the kernel above they superpose into a shared field which every cell samples.
A point in the state space $(\psi_{x}^{*},\psi_{c}^{*})$ was designated as a target morphology, giving prescribed locations and cell types of all cells. This preferred morphology was encoded in a prior and specified desired positions and signal expressions for each clone or cell in terms of the cell’s expectations regarding sensory signals and spatial configurations. This point consisted of the physical form of a head, body, and tail, and with different cell types in each component of the organism. The target sensation under $\lambda$ is plotted in Figure 1(a). In real situations one can imagine a genetic encoding of the target morphology, with epigenetic dynamics being active inference. A possible example of this is given in Figure 1(b) where the target is encoded in binary codons for location and expression. We then equipped each cell with a generative model, specifying it as follows. The mean was given in terms of how environmental states are mapped to sensory states,
@@
with the softmax function returning the expected identity of each cell. In Figure 1(c) we show the softmax function, interpreted as the posterior beliefs that each cell (column) occupies a particular place in the ensemble (rows). A precision $\Pi^{(2)}=e^{-2}$ was also set over external states. A prediction error for how the generative model matched with sensations was given in the form
@@
The model was operationalised by following the procedure in [14]. A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time. A derivative operator $D$ was defined for ordinary differential equations on this space. The entire trajectory was given a coordinate in function space
@@
Now by the decomposition (2.1.1), we have the equations of motion
[…]
for each cell. [^2] This follows from the marginal flow lemma of [35]. This is where the interpretation as inference which we supply as modellers comes into play. We cannot compute the solution to (3.1.1) since the surprisal term yields an intractable integral. However, under the generative model we have equipped our system with, we can instead treat the dynamics of the cell under (3.1.1) as maximising log-evidence for the sensations of the cell, and hence as a variational problem—as posed throughout §2. We will introduce the variational distribution
$),q(\tilde{\psi}\mid\tilde{\mu})=\mathcal{N}\big(\tilde{\mu};-\nabla_{\tilde{\mu}\tilde{\mu}}\log p(\tilde{\mu},\tilde{s},\tilde{a},\tilde{\mu}$$\big)$,
a Gaussian with expectation being the coordinate of the internal trajectory in function space $\tilde{\mu}$ and variance being the curvature of the surprisal—computed at $\tilde{\psi}=\tilde{\mu}$—and free energy
$).F(\tilde{a},\tilde{s},\tilde{\psi})=-\log p(\tilde{s},\tilde{a},\tilde{\mu})+D_{\mathrm{KL}}\big(q(\tilde{\psi}\mid\tilde{\mu})\,\|\,p(\tilde{\psi}\mid\tilde{s},\tilde{a},\tilde{\mu}$$\big)$.
Treating (3.1.1) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final drift terms for the equations of motion for the $i$-th cell given the generative model were obtained as (again denoting the internal state $\psi_{i}$ as $\mu$)
$\begin{split}f_{\tilde{\mu}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})&=(Q_{\mu\mu}-\Gamma_{\mu\mu})\nabla_{\tilde{\mu}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})=D\tilde{\mu}-\nabla_{\tilde{\mu}}\tilde{\varepsilon}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}-\Pi^{(2)}\tilde{\mu}\\ f_{\tilde{a}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})&=(Q_{a_{i}a_{i}}-\Gamma_{a_{i}a_{i}})\nabla_{\tilde{a}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})=D\tilde{a}_{i}-\nabla_{\tilde{a}_{i}}\tilde{s}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}\end{split}$ (3.1.2)
with the following explicit equations for (positional and signalling, respectively) active states
$\partial_{t}\tilde{a}_{x}$ $=D\tilde{a}_{x}-\nabla_{x}\tilde{s}_{x}\cdot\Pi^{(1)}_{x}\tilde{\varepsilon}_{x}+\nabla_{x}\tilde{s}_{\lambda}\cdot\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}$
$\partial_{t}\tilde{a}_{c}$ $=D\tilde{a}_{c}-\Pi^{(1)}_{c}\tilde{\varepsilon}_{c}+\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}.$
As before, if the inference gap is minimised, for instance, in the neighbourhood of the mode, then (3.1.2) is equal to (3.1.1) on the nose.
We then simulated the dynamics of each cell. In Figure 2 we have shown the expected positions and releases of each cell over time, extracted from the solution to (3.1.2), and the decrease in free energy as the cells evolve towards the target state.
In summary, the cell carries a generative model that predicts what signals it should sense if it were the ‘correct’ type of cell at the ‘correct’ place in the target morphology. Those predictions are built from the target locations and target signal combinations, and the model maps internal beliefs about identity through the generative model to predicted sensory consequences. The final outcome of this process is shown in Figure 3.
@@
Figure 1: Plots of the target extracellular gradients (1(a)), encoding of the target signal in the cells (1(b)) and softmax expectations of the identities of each cell (1(c)). In Figure 1(a) colours denote different signal expressions and black dots denote no data, with the final locations of the cells starred.
@@
Figure 2: Here morphogenesis is demonstrated by cells migrating to align with expectations based on sensory input, modelled by a concomitant decrease of free energy in time.
@@
Figure 3: Dynamics of each cell, with final locations starred. The gradients here match those of Figure 2(a). It can be seen that (speaking somewhat heuristically) each cell moves to fulfil its expectations about the signals it should encounter, whilst expressing the signals associated with its current beliefs about its place in the target ensemble.
3.2 Periodically-firing cells
A similar but more complicated example is a ring of excitable, gap-junction-coupled cells whose generative model encodes a periodic target waveform. Each cell’s internal states infer the current phase of that cycle from its membrane sensors, whilst active states like ion-channel gating act to make sensed signals match the predicted periodic trajectory. This model is more expressive because the prior over hidden states is oscillatory, such that the free energy minimum is a limit cycle: the population converges to a stable circulating orbit in state space rather than to a fixed point.
We will define external $\eta_{t}$ as a stimulus in the world, $s_{t}$ as a sensory state, $\mu_{t}$ as the internal state estimating the phase of the cell, and $a_{t}$ as the state of some actuator or motor acting on the environment to make observations conform to predictions. For simplicty we will assume $s_{t}$ tracks $\eta_{t}$ and that sensor noise is standard white in time, i.e. that the likelihood satisfies
[…]
at steady state. We will also introduce a likelihood over outcomes of actions
[…]
in place of the simplifying adiabatic assumption used in the previous example.
The system’s predictions of sensations given a phase is the cosine function $s(\mu)=A\cos\mu$ and we introduce a predictive control prior making actions prefer to create the encoded preference for the waveform,
[…]
There will finally be a uniform prior on $\mu$ as a circle-valued variable.
In total we have a one-step generative model (conditioned on the internal phase):
$s\mid\eta\sim\mathcal{N}(\eta,\sigma_{s}^{2}),\qquad\eta\mid a\sim\mathcal{N}(a,\sigma_{\eta}^{2}),\qquad a\mid\mu\sim\mathcal{N}(A\cos\mu,\sigma_{a}^{2}).$
The generative model is linear in $(s,\eta,a)$ and mildly nonlinear in $\mu$ through the cosine function. To keep this example brief—and to make the surprisal analytically soluble, such that some direct comparisons can be made between (2.1.1) and (2.1.3)—we will not create any hidden states to make inferences about beyond the environmental state, so our generative model is not hierarchical and generalised coordinates of motion need not be used.
In that case, using conditional independence, we can write at steady state (up to additive constants)
$\log p(x)=\frac{(s-\eta)^{2}}{2\sigma_{s}^{2}}+\frac{(\eta-a)^{2}}{2\sigma_{\eta}^{2}}+\frac{(a-A\cos\mu)^{2}}{2\sigma_{a}^{2}}.$
Now specify a parametrisation of $Q$ and $\Gamma$,
$Q=\begin{bmatrix}0&q_{\mu s}&q_{\mu a}&0\\ -q_{\mu s}&0&0&0\\ -q_{\mu a}&0&0&q_{a\eta}\\ 0&0&-q_{a\eta}&0\end{bmatrix},\qquad\qquad\Gamma=\mathrm{diag}(\gamma_{\mu},\gamma_{s},\gamma_{a},\gamma_{\eta}).$
Notice there are no couplings in the flow over the state space between $\mu$ and $\eta$. This is an indication of the Markov blanket.
Since the generative model is somewhat simple, it is straightforward to write the equations of motion:
$\begin{split}\partial_{t}\mu_{t}&=q_{\mu s}\partial_{s}L+q_{\mu a}\partial_{a}L-\gamma_{\mu}\partial_{\mu}L+\xi_{t}^{\mu}\\ &=q_{\mu s}\frac{s-\eta}{\sigma_{s}^{2}}+q_{\mu a}\left(\frac{a-\eta}{\sigma_{\eta}^{2}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)-\gamma_{\mu}\frac{A\sin\mu}{\sigma^{2}_{a}}(a-A\cos\mu)+\xi_{t}^{\mu}\\ \partial_{t}s_{t}&=-q_{\mu s}\partial_{\mu}L-\gamma_{s}\partial_{s}L+\xi_{t}^{s}\\ &=-q_{\mu s}\frac{A\sin\mu}{\sigma^{2}_{a}}(a-A\cos\mu)-\gamma_{s}\frac{s-\eta}{\sigma_{s}^{2}}+\xi_{t}^{s}\\ \partial_{t}a_{t}&=-q_{\mu a}\partial_{\mu}L+q_{a\eta}\partial_{\eta}L-\gamma_{a}\partial_{a}L+\xi_{t}^{a}\\ &=-q_{\mu a}\frac{A\sin\mu}{\sigma^{2}_{a}}(a-A\cos\mu)+q_{a\eta}\left(\frac{\eta-s}{\sigma^{2}_{s}}+\frac{\eta-a}{\sigma_{\eta}^{2}}\right)-\gamma_{a}\left(\frac{a-\eta}{\sigma_{\eta}^{2}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)+\xi_{t}^{a}\\ \partial_{t}\eta_{t}&=-q_{a\eta}\partial_{a}L-\gamma_{\eta}\partial_{\eta}L+\xi_{t}^{\eta}\\ &=-q_{a\eta}\left(\frac{a-\eta}{\sigma_{\eta}^{2}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)-\gamma_{\eta}\left(\frac{\eta-s}{\sigma^{2}_{s}}+\frac{\eta-a}{\sigma_{\eta}^{2}}\right)+\xi_{t}^{\eta}\end{split}$ (3.2.1)
where $\xi^{i}_{t}$ are independent Brownian motions. Notice the solenoidal couplings between sensory states and environmental states preclude any adiabatic assumption.
Using properties of marginals of jointly Gaussian random variables we can compute the steady state
[…]
i.e. i.e. a uniform phase prior, a Gaussian action prior around $A\cos\mu$, and a sensor likelihood concentrated around the action with variance $\sigma_{s}^{2}+\sigma_{\eta}^{2}$. The interpretation of this density is that the triple $(\mu,s,a)$ forms a ribbon-shaped non-equilibrium steady state in $S^{1}\times\mathbf{R}^{2}$ aligned with the subspace $\{a=A\cos\mu,s=a\}$. The solenoidal component makes the probability circulate along this ribbon; however, the shape of the ribbon is fixed by the generative terms specified above.
The system can treated as a linear Gaussian chain, with Gaussian likelihoods for $s\mid\eta$ and $\eta\mid s,a$, providing an exact expression for the posterior
[…]
Completing the square in $\eta$ yields $p(\eta\mid s,a)=\mathcal{N}(m_{\mathrm{post}},v_{\mathrm{post}})$ where
$v_{\mathrm{post}}=\left(\frac{1}{\sigma_{s}^{2}}+\frac{1}{\sigma_{\eta}^{2}}\right)^{-1},\qquad m_{\mathrm{post}}=v_{\mathrm{post}}\left(\frac{s}{\sigma^{2}_{s}}+\frac{a}{\sigma^{2}_{\eta}}\right)$
so that the posterior is parametrised by a variance of twice the inverse precision. Now let the internal states parametrise a Gaussian posterior expressing beliefs about the world, $q(\eta;\mu,\tau^{2})=\mathcal{N}(\eta;A\cos\mu,\tau^{2})$. The free energy can be expressed analytically as
$F(s,a,\mu)$ $=\operatorname{\mathbb{E}}_{q}[\log q(\eta;\mu,\tau^{2})]-\operatorname{\mathbb{E}}_{q}[\log p(\eta\mid s,a)]-\log p(s,a,\mu)$
$=-\frac{1}{2}\log 2\pi\tau^{2}-\frac{1}{2}+\frac{1}{2}\log 2\pi v_{\mathrm{post}}+\frac{1}{2v_{\mathrm{post}}}\left(\tau^{2}+(A\cos\mu-m_{\mathrm{post}})^{2}\right)-\log p(s,a,\mu)$
or, collecting like terms,
$\frac{1}{2}\left(\log\frac{v_{\mathrm{post}}}{\tau^{2}}+\frac{\tau^{2}+(A\cos\mu-m_{\mathrm{post}})^{2}}{v_{\mathrm{post}}}-1\right)-\log p(s,a,\mu)$
so that when $\tau^{2}=v_{\mathrm{post}}$ and $m_{\mathrm{post}}=A\cos\mu$, the KL term vanishes.
Taking gradients of the free energy yields the equations of motion
$\begin{split}\partial_{t}\mu_{t}&=q_{\mu s}\partial_{s}F+q_{\mu a}\partial_{a}F-\gamma_{\mu}\partial_{\mu}F+\xi_{t}^{\mu}\\ &=q_{\mu s}\frac{s-A\cos\mu}{\sigma_{s}^{2}}+q_{\mu a}(a-A\cos\mu)\left(\frac{1}{\sigma_{\eta}^{2}}+\frac{1}{\sigma_{a}^{2}}\right)\\ &\phantom{--}-\gamma_{\mu}A\sin\mu\left(\frac{s-A\cos\mu}{\sigma_{s}^{2}}+\frac{a-A\cos\mu}{\sigma^{2}_{\eta}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)+\xi_{t}^{\mu}\\ \partial_{t}a_{t}&=-q_{\mu a}\partial_{\mu}F-\gamma_{a}\partial_{a}F+\xi_{t}^{a}\\ &=-q_{\mu a}A\sin\mu\left(\frac{s-A\cos\mu}{\sigma_{s}^{2}}+\frac{a-A\cos\mu}{\sigma^{2}_{\eta}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)\\ &\phantom{--}-\gamma_{a}(a-A\cos\mu)\left(\frac{1}{\sigma_{\eta}^{2}}+\frac{1}{\sigma_{a}^{2}}\right)+\xi_{t}^{a}.\end{split}$ (3.2.2)
We then simulated the system as a Kalman update for a linear Gaussian state space channel $\eta\to s$ with input aa. [^3] The python code for this simulation is available at https://github.com/DARSakthi/periodically_firing_cells/. In practice, this means when simulating the system forwards, we set the parameters of the variational density to those of the posterior such that the variational density $q$ became the (instantaneous) Kalman posterior over the latent world state $\eta$ and the instantaneous precision-weighted prediction errors drove the gradient descent on $F$.
In Figure 4 sampled trajectories of $\mu_{t}$ under (3.2.1) and (3.2.2) are compared (Figure 5, $a_{t}$, respectively). Close correlation is observed between the two flow regimes. A corresponding tendency towards zero of the sample mean of the KL term is observed in Figure 6. Plotted per-timepoint sample means of the free energy under both flow regimes are contained in Figure 7. By adding the precision-weighted prediction error (KL) term to the internal updates, $F$ has a faster decrease and a greater magnitude of decrease under (3.2.2); this is evidently from the prediction error driving the flow of $\partial_{\mu}F$ to align $q(\eta;\mu)$ with $p(\eta\mid s,a)$.
Figure 4: Evolution of the internal state under $\mathcal{L}$ and under $F$. The dynamics of internal states in two regimes, plotted on a polar spiral showing the evolution in the phase in time. The dynamics under a gradient descent on $\mathcal{L}$ and on $F$ show excellent agreement over all timescales.
Figure 5: Evolution of the active state under $\mathcal{L}$ and under $F$. The dynamics of the active state in two regimes were also plotted; again, excellent agreement is observed.
Figure 6: Decrease of the KL divergence under the flow on $F$. The sample mean ($N=200$) of the KL divergence along trajectories of (3.2.2) is shown.
Figure 7: Decrease of the free energy under trajectories in both flow regimes. The sample mean ($N=200$) of the variational free energy along trajectories of (3.2.1) and along trajectories of (3.2.2) are compared. Both drive the free energy towards zero, however, the flow under $F$ decreases more rapidly and attains a lesser minimum closer to zero.
The reason an inferential interpretation is useful here is in the coding of stimuli in neurones and neuronal ensembles. Minimising variational free energy, and the resultant model of internal and active states behaving as if they were performing Bayesian inference, endows the phase of the oscillation with the interpretation of a posterior belief about expected incoming signals in the environment and the Kalman-like updates as modulations of gain to minimise precision-weighted prediction error, rather than merely a position on a limit cycle. Under the relationship to stochastic thermodynamics forged in e.g. [24], quantitative predictions can be made about the cost of maintaining such beliefs—and with an explicit low-dimensional representation of such signals given by the concentration of the non-equilibrium steady state density, we have an organically emerging neural manifold from those beliefs.